# Decoding Safety Feedback from Diverse Raters: A Data-driven Lens on Responsiveness to Severity

### v4

@@ Abstract
Ensuring the safety of Generative AI requires a nuanced understanding of pluralistic viewpoints. In this paper, we introduce a novel data-driven approach for interpreting granular ratings in pluralistic datasets. Specifically, we address the challenge of analyzing nuanced differences in safety feedback from a diverse population expressed via ordinal scales (e.g., a Likert scale). We distill non-parametric responsiveness metrics that quantify the consistency of raters in scoring varying levels of the severity of safety violations. Leveraging a publicly available pluralistic dataset of safety feedback on AI-generated content as our case study, we investigate how raters from different demographic groups (age, gender, ethnicity) use an ordinal scale to express their perceptions of the severity of violations. We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems reliably in multi-cultural contexts. We show that our approach can inform rater selection and feedback interpretation by capturing nuanced viewpoints across different demographic groups, hence improving the quality of pluralistic data collection and in turn contributing to more robust AI development.
@@ 1 Introduction
Ensuring the safety of Generative AI is paramount for their responsible deployment and societal trust. Recent research demonstrates that safety perceptions are not uniform but vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024).
Safety evaluation tasks often use binary ratings, such as safe and unsafe, which lack the granularity needed for effective alignment with human preferences (Wu et al., 2023; Collins et al., 2024). To capture more fine-grained perceptions of the severity of potential harm, safety alignment tasks increasingly employ ordinal scales, such as Likert scales (Curry et al., 2021), which allow for more nuanced feedback that is critical for the creation of pluralistic datasets for SFT / RLHF. However, these scales are harder to interpret due to variations in individual perceptions and response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales). Inconsistencies in how raters map perceived severity to ordinal scores can propagate to datasets and optimization objectives for safety alignment, especially in multi-cultural contexts (Kaufmann et al., 2024). This, in turn, can lead to exaggerated safety behaviors in models like LLMs, e.g., false refusal of benign requests or generation of content that is harmful to certain cultures. Moreover, when LLMs are used as judges or reward models (Bavaresco et al., 2024; Wang et al., 2024), such inconsistencies and response biases can degrade their utility. Both cases cause downstream harms due to poor understanding of nuanced safety feedback.
This paper introduces a novel data-driven non-parametric approach for interpreting granular ratings in pluralistic datasets, offering more robust and nuanced insights than traditional approaches. Using safety feedback on AI-generated content as a case study, we address the critical challenge of understanding how raters from diverse demographic groups utilize ordinal scales (e.g., Likert scales) when expressing the severity of safety violations. The main contributions of this paper are:
Metrics development: We distill robust responsiveness metrics from observable data to analyze the feedback of individual raters or rater groups for varying levels of severity. These metrics allow us to:
@@
Metrics application: We apply these metrics to a publicly available pluralistic dataset of safety feedback on AI-generated content (Rastogi et al., 2024), demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems in multi-cultural contexts, specifically:
  • Understanding scale usage: uncovering patterns in the use of the given Likert scale, thus elucidating genuine variations in the expressions of demographic groups;
  • Capturing nuanced viewpoints: quantifying the responsiveness of demographic groups to severity across violation types, resulting in a granular understanding of pluralistic viewpoints;
  • Prioritizing high-impact items: taking items deemed highly unsafe by raters with high responsiveness to severity from different groups;
  • Improving pluralistic data collection: establishing a reliable and repeatable process for selecting raters with high responsiveness to severity from different groups.
@@ 2 Related work
Our work builds on the state-of-the-art in eliciting nuanced human feedback in Generative AI evaluation and expands existing research on calibration of human judgments.
@@
Nuanced human feedback for AI safety. Collecting human perspectives on behavior of generative AI models is exceedingly commonplace with growth in its usage in real world tasks. Across the literature on AI evaluation, different configurations of human feedback have been studied, with an outsized focus on binary (0/1) human feedback. Recent research (Collins et al., 2024; Arhin et al., 2021; Denton et al., 2021) discusses the limitations of binary feedback in capturing the nuance involved in generative AI evaluation, especially in safety. Wu et al. (2023); Collins et al. (2024) propose fine-grained human feedback encompassing evaluation across multiple attributes and with higher density, yielding improvement in downstream AI tasks via RLHF. Further, Rauh et al. (2024); Jiang et al. (2021) emphasize the importance of measuring extent of harm (severity) in evaluation of algorithms. Another dimension in collecting human feedback relates to the identity of the human providing the feedback. The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023; Pei & Jurgens, 2023; Davani et al., 2024). For developing AI that aligns with human values, Sorensen et al. (2024) show the importance of considering pluralistic viewpoints from a diverse set of raters. Our research builds upon this body of work by specifically examining human feedback collected on a fine-grained ordinal scale from raters belonging to different groups with different collective identities.
@@
Interpretation and calibration of human judgments. Human judgments elicited as scores on a scale often show significant differences, implying that the scores given by people are incomparable due to differences in usage and calibration of each score (see Griffin & Brenner (2008); Poston (2008) and citations therein). Such differences in human scores are usually interpreted through simplifying modeling assumptions about how miscalibrations present in the data. These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner & Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang & Shah, 2019). However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin & Brenner (2008) and citations therein). In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to quantify the consistency of raters in reflecting varying levels of severity. Traditional non-parametric metrics like Kendall’s $\tau$ and area under the PR or ROC curves do not capture well the responsiveness to severity. Using a real-world dataset, we surface insights from our proposed metrics into the feedback patterns of different rater groups.
@@ 3 Setup
    Trained raters strictly follow a set of prescribed rules. For each item, they provide a binary score of 0 or 1 indicating their feedback of safe or unsafe respectively.
@@ 3.1 Data model
where $f_{j}$ is the rater-specific perception function. This model acknowledges that true underlying severity $V$ of items is not directly measurable but is still the principal factor that monotonically influences every rater’s judgment. In the case of trained raters, $f_{j}$ represents the shared understanding of severity prescribed by their strict guidelines that serve as a framework to operationalize the theoretical notion of severity. In the case of crowd raters, $f_{j}$ represents some perception of true severity based on their lived experiences. In other words, perception of rater $j$ is primarily distributed along the dimension $V^{j}$ such that $V^{j}_{i}$ represents the unobservable true severity of item $i$ as perceived by rater $j$. Hereon, we use the notation $V^{j}$ for true severity as perceived by rater $j$ and the notation $V$ for the true underlying severity, aka simply, true severity.
@@
We assume $V$ to be unidimensional following the standard assumption of a unidimensional latent trait found in methods like Item Response Theory (Samejima, 1968). We note that the assumption limits our ability to simultaneously measure responsiveness with respect to multiple dimensions of severity.
@@ 3.2 Responsiveness to severity
Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to convey both granular and broader variations in the true underlying severity $V$ using the Likert scale. Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum & Navarrete, 1998). The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot & Fienberg, 1983) in the case that we consider calibration of crowd rater scores against some binarized reference for severity. The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivate the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938). Furthermore, the property is also related to the notion of discrimination between different levels of the latent trait in Item Response Theory (Samejima, 1968) and Mokken Scale Analysis (Mokken, 1971).
@@ 4 Metric design
Next, our aim is to quantify the responsiveness of the diverse crowd raters to the severity of violations in order to evaluate and compare them. Such a quantification can be achieved by individually quantifying the two properties that constitute responsiveness. We note that during safety alignment, the notion of true severity $V$ is operationalized depending on the choice of the feedback paradigm. When safety alignment is conducted based on feedback of trained raters, true severity $V$ is operationalized as captured by their guidelines, i.e., $V^{g}$. When safety alignment is conducted based on feedback of crowd raters, true severity $V$ is operationalized as captured by the collective judgment of the crowd, i.e., $V^{c}$. In the latter case, there is no dependence on the existence of any guidelines.
Either way, true severity $V$ remains directly unobservable. We take $U$ to denote the observable binary reference such that $U=1$ signals $V>T$, i.e., unsafe. We can now reformulate the inequalities presented above for the ability to stochastically order and the ability to discriminate by leveraging $U=1$ in place of $V>T$ and $U=0$ in place of $V\leq T$. Note that since we do not assume a fixed threshold $T$, it is not straightforward to directly substitute $U=1$ into the threshold-specific inequalities. However, we show that if the two inequalities hold for all individual thresholds $T$, then they also hold for $U=1$. Proofs for both the inequalities are in appendix A.
  • Ability to stochastically order. We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ when $s_{1}>s_{2}$, for all $T=t$. Hence, we can prove that $s_{1}>s_{2}$, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
@@
Depending on the operationalization of true severity $V$ under consideration, we can obtain the observable binary reference $U$, i.e., guideline-based reference or crowd-based reference, as described:
@@
    We can treat the binary scores of the trained raters as the $U$. We assign the individual binary scores of every trained rater to the items by replicating each item for every trained rater. When $U$ is obtained from trained raters, we are quantifying the responsiveness to severity $V^{g}$.
@@
    We can derive the $U$ from the crowd raters by excluding the rater(s) being evaluated to maintain the independence of $S$ and $T$ by construction. We binarize the individual Likert scores of the crowd raters, excluding those being evaluated, using scores $k\in[1,K]$ as boundaries. We assign these individual binary scores to the items by replicating each item for every crowd rater. When $U$ is obtained from crowd raters, we are quantifying the responsiveness to severity $V^{c}$.
The design of our metrics itself is agnostic to the choice of the reference, and we work with both.
@@ 4.1 Metrics for the two properties
We formulate metrics for the two properties based on the standard concepts of precision and recall. Given a Likert scale with scores $S\in\{0,1,2,3,\dots,K\}$, we take $Precision(S)$ to denote the precision when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe. Similarly, we take $Recall(S)$ to denote the recall when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe. The decision to compute precisions and recalls exactly at $S$ rather than $\geq S$ is a crucial one to our metric development because we directly leverage the terms $P(U=1|S=s)$ and $P(S=s|U=1)$. We define the following metrics to quantify the strength of the core inequalities for the two properties:
Monotonic Precision Area for stochastic ordering. We note that the probability $P(U=1|S=s)$ is equivalent to $Precision(s)$, i.e., the precision when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe. Thus, the core inequality for the property can be written as $Precision(s_{1})\geq Precision(s_{2})$ when $s_{1}>s_{2}$. We take the area under the curve defined by
@@
at $s=0,1,2,3,\dots,K$, to directly and non-parametrically quantify increases in $P(U=1|S=s)$ while penalizing any violations of monotonicity at each score. Since $0\leq Precision(\cdot)\leq 1$, it can be shown via linear programming that the maximum possible area is given by $\left\lceil\frac{K+1}{2}\right\rceil*\left\lfloor\frac{K+1}{2}\right\rfloor$. To explain it intuitively, the maximum possible area is achieved when $Precision(s)$ is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater has perfectly aligned the midpoint of the Likert scale with the threshold separating $U=1$ (unsafe) and $U=0$ (safe). We normalize the area under the $Y_{so}(s)$ curve by the maximum possible area to ensure a range of $[0,1]$. Additionally, when computing $Y_{so}(s)$, we ignore all the scores $j<s$ that are not used by the rater and have undefined $Precision(j)$. Similarly, if score $s$ is not used by the rater, we take $Y_{so}(s)=0$.
Weighted Recall Area for discrimination. We note that the probability $P(S=s|U=1)$ is equivalent to $Recall(s)$, i.e., the recall when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe. Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall when classifying items with score $S<s$ as safe and items with score $S\geq s$ as unsafe. Since $\sum Recall(\cdot)=1$, without needing normalization, we take the area under the curve defined by
@@
at $s=0,1,2,3,\dots,K$. A high area means high $P(S<s|U=0)$ and $P(S=s|U=1)$, which directly implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property. Essentially, $Y_{d}(s)$ is the recall of unsafe items at score $s$, i.e., $Recall(s)$, weighted by the proportion of safe items correctly discriminated by scores $<s$. It represents the concordance probability (Heller & Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $S=s$ to unsafe items while assigning scores $S<s$ to safe items.
Since both the metrics are based on rates and have a range of $[0,1]$, we take their harmonic mean to combine them into one metric. We have four cases to consider here. When a rater $j$ exhibits high MPA and high WRA, their $V^{j}$ is aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\approx V$, and they are responsive to both granular variations as well as broader distinctions in $V$. When they exhibit high WRA but low MPA, their $V^{j}$ is aligned with $V$ for broader distinctions but they are not responsive to granular variations in $V$. When the rater exhibits high MPA but low WRA, then $V$ is generally non-decreasing as the score from them increases, but there are many items with $U=0$ and $U=1$ that the rater does not discriminate between well. Finally, when the rater exhibits low MPA and low WRA, their $V^{j}$ is not aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\neq V$. We note that both the metrics are provably statistically consistent. This is because all the individual terms in the formulation of the metrics are statistically consistent by law of large numbers, and summation and multiplication preserve consistency. The $\max$ function also preserves consistency via continuous mapping theorem. Appendix B provides an implementation of our metrics in code.
4.2 Contrasting with traditional metrics
There are potentially many alternative non-parametric approaches that could be used to assess responsiveness to severity. For example, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve (AUROC), or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether or not the relationship between crowd rater scores and the reference is monotonic. We highlight some limitations of these approaches below:

### v5

@@ Abstract
Ensuring the safety of Generative AI requires a nuanced understanding of pluralistic viewpoints. In this paper, we introduce a novel data-driven approach for analyzing ordinal safety ratings in pluralistic settings. Specifically, we address the challenge of interpreting nuanced differences in safety feedback from a diverse population expressed via ordinal scales (e.g., a Likert scale). We define non-parametric responsiveness metrics that quantify how raters convey broader distinctions and granular variations in the severity of safety violations. Leveraging publicly available datasets of pluralistic safety feedback as our case studies, we investigate how raters from different demographic groups use an ordinal scale to express their perceptions of the severity of violations. We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems reliably in multi-cultural contexts. We show that our approach can inform rater selection and feedback interpretation by capturing nuanced viewpoints across different demographic groups, hence improving the quality of pluralistic data collection and in turn contributing to more robust AI alignment.
@@ 1 Introduction
Ensuring the safety of Generative AI is paramount for their responsible deployment and societal trust. Safety evaluation tasks that use binary ratings, such as safe and unsafe, lack the granularity needed for effective alignment with human preferences (Wu et al., 2023; Collins et al., 2024). To capture more fine-grained perceptions of the severity of potential harm, safety alignment tasks increasingly employ ordinal scales, such as Likert scales (Curry et al., 2021), which allow for more nuanced feedback that is vital to the creation of alignment datasets for Supervised Fine-Tuning (SFT) / Reinforcement Learning from Human Feedback (RLHF). A significant challenge, however, is that these scales are harder to interpret due to variations in individual perceptions and response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales). Such biases can propagate to datasets and optimization objectives for safety alignment (Kaufmann et al., 2024), leading to exaggerated safety behaviors in Generative AI, e.g., false refusal of benign requests.
Yet another significant challenge that emerges from recent research is that safety perceptions are not uniform; they vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024). Poor understanding of differences in how diverse raters or rater groups perceive the severity of violations and map the perceived severity to ordinal scores can lead to the design of optimization objectives that are insensitive or even harmful to certain cultures. Moreover, when LLMs trained on such pluralistic ordinal ratings are used as judges or reward models (Bavaresco et al., 2024; Wang et al., 2024), they may not represent well the granular variations or even broader distinctions in safety perceptions of different raters or rater groups. Both cases cause downstream harms due to poor understanding of nuanced safety feedback.
This paper introduces a novel data-driven non-parametric approach for analyzing nuanced pluralistic safety feedback, offering deeper and more robust insights than traditional non-parametric and parametric approaches. Leveraging two publicly available datasets of pluralistic safety feedback as our case studies, we address the challenge of understanding the differences in ordinal ratings from diverse demographic groups when expressing the severity of safety violations. The main contributions of this paper are:
Metrics development: We define robust responsiveness metrics from observable data to analyze the feedback of individual raters or rater groups for varying levels of severity. These metrics allow us to:
@@
Metrics application: We apply these metrics to two publicly available pluralistic datasets of safety feedback, demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems in multi-cultural contexts, specifically:
  • Understanding scale usage: uncovering patterns in the use of the given Likert scale, thus elucidating pluralistic variations in the expressions of demographic groups;
  • Capturing nuanced viewpoints: quantifying the responsiveness of demographic groups to severity across violation types, resulting in a deeper understanding of pluralistic viewpoints;
  • Informing pluralistic data collection: establishing a reliable and repeatable process for selecting raters with high responsiveness to severity from different groups;
  • Guiding safety alignment in pluralistic settings: reflecting the granular variations in safety perceptions of diverse raters alongside the broader distinctions in their safety perceptions.
@@ 2 Related work
Our work builds on the state-of-the-art in eliciting nuanced human feedback in Generative AI evaluation and expands existing research on understanding of human judgments.
@@
Nuanced human feedback for AI safety. Collecting human perspectives on behavior of generative AI models is exceedingly commonplace with growth in its usage in real world tasks. Across the literature on AI evaluation, different configurations of human feedback have been studied, with an outsized focus on binary (0/1) human feedback. Recent research (Collins et al., 2024; Arhin et al., 2021; Denton et al., 2021) discusses the limitations of binary feedback in capturing the nuance involved in generative AI evaluation, especially in safety. Wu et al. (2023); Collins et al. (2024) propose fine-grained human feedback encompassing evaluation across multiple attributes and with higher density, yielding improvement in downstream AI tasks via RLHF. Further, Rauh et al. (2024); Jiang et al. (2021) emphasize the importance of measuring extent of harm (severity) in evaluation of algorithms. Another dimension in collecting human feedback relates to the identity of the human providing the feedback. The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023; Pei and Jurgens, 2023; Davani et al., 2024). For developing AI that aligns with human values, Sorensen et al. (2024) show the importance of considering pluralistic viewpoints from a diverse set of raters. Our research builds upon this body of work by specifically examining human feedback collected on a fine-grained ordinal scale from raters belonging to different groups with different collective identities.
@@
Interpretation and calibration of human judgments. Human judgments elicited as scores on a scale often show significant differences, implying that the scores given by people are incomparable due to differences in usage and calibration of each score (see Griffin and Brenner (2008); Poston (2008) and citations therein). Such differences in human scores are usually interpreted through simplifying modeling assumptions about how miscalibrations present in the data. These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner and Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang and Shah, 2019). However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin and Brenner (2008) and citations therein). In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to quantify the consistency of raters in reflecting varying levels of severity. Traditional non-parametric metrics like Kendall’s $\tau$ and area under the PR or ROC curves do not capture well the responsiveness to severity. Using real-world datasets, we surface insights from our proposed metrics into the feedback patterns of different rater groups.
@@ 3 Setup
    Trained raters strictly follow a set of prescribed guidelines. For each item, they provide a binary score of 0 or 1 indicating their feedback of safe or unsafe respectively.
@@ 3.1 Data model
where $f_{j}$ is the rater-specific perception function. This model acknowledges that true underlying severity $V$ of items is not directly measurable but is still the principal factor that monotonically influences every rater’s judgment. In the case of trained raters, $f_{j}$ represents the shared understanding of severity prescribed by their strict guidelines that serve as a framework to operationalize the theoretical notion of severity. In the case of crowd raters, $f_{j}$ represents some perception of true severity based on their lived experiences. In other words, perception of rater $j$ is primarily distributed along the dimension $V^{j}$ such that $V^{j}_{i}$ represents the unobservable true severity of item $i$ as perceived by rater $j$. Hereon, we use the notation $V^{j}$ for true severity as perceived by rater $j$ and the notation $V$ for the true underlying severity, or simply, true severity.
@@
We assume $V$ to be unidimensional following the standard assumption of a unidimensional latent trait found in frameworks like Item Response Theory (Samejima, 1968). We note that if $V$ is taken to be multidimensional, then our metrics can be applied per dimension provided that scores from the raters are also collected per dimension. But more importantly, unlike Item Response Theory (IRT), we do not assume some common latent calibrated scale(s) across all the raters, i.e., $V^{j}$ are specific to the raters, and we impose no further restrictions on them. This is crucial for pluralistic settings.
@@ 3.2 Responsiveness to severity
Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to convey both granular and broader variations in the true underlying severity $V$ using the Likert scale. Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum and Navarrete, 1998). The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot and Fienberg, 1983) in the case that we consider calibration of crowd rater scores against some binarized reference for severity. The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivate the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938). Furthermore, the property is also related to the notion of discrimination between different levels of the latent trait in Item Response Theory (Samejima, 1968) and Mokken Scale Analysis (Mokken, 1971).
@@ 4 Metric design
Next, our aim is to quantify the responsiveness of crowd raters using Likert scales to the severity of violations in order to evaluate and compare them. Such a quantification can be achieved by individually quantifying the two properties that constitute responsiveness.
4.1 Reference for true severity
We note that during safety alignment, the notion of true severity $V$ is operationalized depending on the choice of the feedback paradigm. When safety alignment is conducted based on the feedback of trained raters, true severity $V$ is operationalized as captured by their guidelines, i.e., $V^{g}$. When safety alignment is conducted based on the feedback of crowd raters, true severity $V$ is operationalized as captured by the collective judgment of the crowd, i.e., $V^{c}$. In the latter case, there is no dependence on the existence of any guidelines. There can be other possible operationalizations too, but we focus on these two.
In any operationalization, true severity $V$ remains directly unobservable. We take $U$ to denote the observable binary reference such that $U=1$ signals $V>T$ and $U=0$ signals $V\leq T$. We can now reformulate the inequalities presented above for the ability to stochastically order and the ability to discriminate by leveraging $U=1$ in place of $V>T$ and $U=0$ in place of $V\leq T$. Note that since we do not assume a fixed threshold $T$, it is not straightforward to directly substitute $U=1$ into the threshold-specific inequalities. However, we show that if the two inequalities hold for all individual thresholds $T$, then they also hold for $U=1$. Proofs for both the inequalities are in appendix A.
  • Ability to stochastically order. We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ when $s_{1}>s_{2}$, for all $T=t$. Hence, we can prove that $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$ when $s_{1}>s_{2}$.
@@ 4.1 Reference for true severity
Depending on the two operationalizations of true severity $V$ under consideration, we can obtain the observable binary reference $U$, i.e., guideline-based reference or crowd-based reference, as described:
@@
    We can treat the binary scores of trained raters as the binary reference $U$. We assign the individual binary scores of trained raters directly to the respective items. Here, $S$ and $T$ are trivially independent. When $U$ is obtained from trained raters, we are quantifying the responsiveness to severity $V^{g}$.
@@
    We can derive the binary reference $U$ from crowd raters themselves, excluding the rater(s) being evaluated to maintain non-circularity. Using points 1 to $K$ on the scale as exogenous boundaries to maintain the independence of $S$ and $T$, we obtain the binary reference $U$ per boundary by binarizing the individual Likert scores of crowd raters, excluding those being evaluated. When $U$ is obtained from crowd raters themselves, we are quantifying the responsiveness to severity $V^{c}$.
The design of our metrics itself is agnostic to the choice of the reference, and we work with both. References can have their own properties, so responsiveness must be interpreted in light of the chosen operationalization.
@@ 4.2 Metrics for the two properties
We formulate metrics for the two properties based on the standard concepts of precision and recall. Given a Likert scale with scores $S\in\{0,1,2,3,\dots,K\}$, $Precision(s)$ denotes the precision when score $S=s$ is taken as positive and all scores $S\neq s$ are taken as negative. Similarly, $Recall(s)$ denotes the recall when score $S=s$ is taken as positive and all scores $S\neq s$ are taken as negative. The decision to compute precisions and recalls exactly at $S=s$ rather than $S\geq s$ is crucial to our metric development because we directly leverage the terms $P(U=1|S=s)$ and $P(S=s|U=1)$. We define the following metrics to quantify the strength of the core inequalities for the two properties:
Monotonic Precision Area for stochastic ordering. We noted that the probability $P(U=1|S=s)$ is equivalent to $Precision(s)$, i.e., the precision when taking items with score $S=s$ to be the positives and items with scores $S\neq s$ to be the negatives. Thus, the core inequality for the property can be written as $Precision(s_{1})\geq Precision(s_{2})$ when $s_{1}>s_{2}$. We take the area under the curve defined by
@@
at $s=0,1,2,3,\dots,K$, to directly and non-parametrically quantify increases in $P(U=1|S=s)$ while penalizing any violations of monotonicity at each score. Since $0\leq Precision(\cdot)\leq 1$, it can be shown via linear programming that the maximum possible area is given by $\left\lceil\frac{K+1}{2}\right\rceil*\left\lfloor\frac{K+1}{2}\right\rfloor$. To explain it intuitively, the maximum possible area is achieved when $Precision(s)$ is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater uses the Likert scale in a perfectly balanced way while also having no violations of monotonicity. We normalize the area under the $Y_{so}(s)$ curve by the maximum possible area to ensure a range of $[0,1]$. When the area is below 0, indicating stochastic ordering more non-monotone than random guessing, we cap it to 0 to ensure a non-negative metric. Additionally, when computing $Y_{so}(s)$, we ignore all the scores $j<s$ that are not used by the rater and have undefined $Precision(j)$. Similarly, if score $s$ is not used by the rater, we take $Y_{so}(s)=0$.
Weighted Recall Area for discrimination. We noted that the probability $P(S=s|U=1)$ is equivalent to $Recall(s)$, i.e., the recall of items with $U=1$ when score $S=s$ is taken as positive and scores $S\neq s$ as negative. Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall of items with $U=0$ when scores $S<s$ are taken as positive and scores $S\geq s$ as negative. Since $\sum Recall(\cdot)=1$, without needing normalization, we take the area under the curve defined by
@@
at $s=0,1,2,3,\dots,K$. A high area means high $P(S<s|U=0)$ and $P(S=s|U=1)$, which directly implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property. Essentially, $Y_{d}(s)$ is the recall of items with $U=1$ at score $S=s$ weighted by the recall of items with $U=0$ discriminated by scores $S<s$. It represents the concordance probability (Heller and Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $S=s$ to items with $U=1$ while assigning scores $S<s$ to items with $U=0$.
Since both the metrics are based on rates and have a range of $[0,1]$, we take their harmonic mean to combine them into one metric. We have four scenarios to consider here. When a rater $j$ exhibits high MPA and high WRA, their $V^{j}$ is aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\approx V$, and they are responsive to both granular variations as well as broader distinctions in $V$. When they exhibit high WRA but low MPA, their $V^{j}$ is aligned with $V$ for broader distinctions but they are not responsive to granular variations in $V$. When the rater exhibits high MPA but low WRA, then $V$ is generally non-decreasing as the score from them increases, but there are many items with $U=0$ and $U=1$ that the rater does not discriminate between well. Finally, when the rater exhibits low MPA and low WRA, their $V^{j}$ is not aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\neq V$. In appendix B, we provide visualizations and recommendations for the different scenarios.
We note that both the metrics are provably statistically consistent. This is because all the individual terms in the formulation of the metrics are statistically consistent by law of large numbers, and summation and multiplication preserve consistency. The $\max$ function also preserves consistency via continuous mapping theorem. Appendix C provides an implementation of our metrics in code.
4.3 Contrasting with other metrics
There are potentially many other parametric and non-parametric approaches that could be used to assess responsiveness to severity. Looking at non-parametric options, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve (AUROC), or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether or not the relationship between crowd rater scores and the reference is monotonic. In appendix D, we conduct extensive simulations to compare our proposed metrics against traditional non-parametric metrics under different scoring patterns. We highlight some limitations of the traditional metrics that emerge: